OpenStudy (anonymous):

solve by substitution 1/3x+1/5y=-4/5 1/3x+2y=10

OpenStudy (anonymous):

Simplest is to note that the second equation converts to 1/3x = 10- 2y and replace 1/3x in the first equation. Then solve for y and compute x.

Directrix (directrix):

(1/3)x+(1/5)y=-4/5 (1/3) x+2y=10 -------------- (1/3) x =10 - 2y 10 - 2y + (1/5)y=-4/5 after subsitution See if you get x = -6 and y = 6.